# Awes, Laws, and Flaws From Today’s LLM Research

### v2

@@ Abstract
We perform a critical examination of the scientific methodology behind contemporary large language model (LLM) research. For this we assess over 2,000 research works based on criteria typical of what is considered good research (e.g. presence of statistical tests and reproducibility) and cross-validate it with arguments that are at the centre of controversy (e.g., claims of emergent behaviour, the use of LLMs as evaluators). We find multiple trends, such as declines in claims of emergent behaviour and ethics disclaimers; the rise of LLMs as evaluators in spite of a lack of consensus from the community about their useability; and an increase of claims of LLM reasoning abilities, typically without leveraging human evaluation. This paper underscores the need for more scrutiny and rigour by and from this field to live up to the fundamentals of a responsible scientific method that is ethical, reproducible, systematic, and open to criticism.
@@ 1 Introduction
Large language models (LLMs) [^1] We use the term “LLM” loosely to refer to generative text-to-text models with sizes above 1B parameters. are a powerful technology. They can follow instructions and output coherent, persuasive text. This has made them the centre of attention in academia, industry, and the media. Given the potential funding associated with these technologies, it should not be a surprise that news and research works are sometimes accompanied by bold claims about their capabilities.
However, it is known that some areas of AI research have issues with reproduciblity and other experimentation protocols. For example, research works do not provide sufficient details for independent verification [1, 2]; or they report aggregate performances (e.g., accuracy) without providing detailed protocols or error breakdowns [3, 1, 4, 5, 2]. This is especially common in health sciences [6], security [7], and recommender systems [8]–all areas where LLMs are increasingly being used. Even though studies and meta-reviews usually come with suggested good practices (reporting carbon footprints [9]; reproducibility checklists and experimental protocols [10, 11, 12]; self-contained artifacts [13]; and metadata for corpora [14]), these suggestions are often not heeded. Last year, Gehrmann et al. [1] recommended good practices for research, and noted that, out of 66 articles from leading conferences, only between 15% and 35% of the recommendations were partially followed. The practices themselves might not even be sufficiently impactful [7], given that other experimental protocols (e.g., sampling, initialisation, hyperparameters) may also impact reproducibility [3]. These are issues particularly relevant to LLMs, as they are intrinsically stochastic and sometimes limited to an API call.
@@
Given all this, it has been said that the focus of AI research is the approaches, rather than the results, or, rather, that less importance is allocated to issues around experimental protocols when compared to other fields [4]. We argue, however, that in LLM research there has been a shift towards result-driven experimentation: partly due to their high availability and remarkable capabilities, but also due to the aforementioned combination of capabilities and attention from stakeholders. This in turn suggests that the scientific community should allocate the same relevance to experimental protocols and good research practices as in other fields.
@@
Indeed, LLM research is no stranger to reproducibility and documentation problems. Reviews have found that the most downloaded models in Huggingface do not consistently provide the same amount of information in their documentation [15]; and a recent analysis on the experimental protocols of over 40 chat-based LLMs (e.g., availability of data, weights, licences, etc) found that, while many projects claimed some level of open-sourcing, documentation was “exceedingly rare” [5].
What makes LLM research particularly difficult, however, goes beyond self-reporting. The technology itself has challenges, such as the closed-source nature of some models, the high cost of training the larger versions, and that reproducibility could be limited to an API call with adjustable parameters. To add, they are increasingly treated like software, with major and minor versions; as well as “stealth” updates to the same model. Conversely, prompting LLMs to solve a task is extremely easy. Their availability, coupled with their functionality as writing assistants, could lead to a rapidly-increasing volume of scientific articles with strong claims and lacklustre experimental practices.
This is not a criticism: evaluation of these models is notoriously difficult, and researchers are under pressure to keep up. Scaling this evaluation is challenging due to the length and complexity of their output, combined with the rapid turnaround of this technology. Indeed, automated metrics like BLEU are known to not capture natural language generation well [16, 17], and not correlate well with human judgements [18, 1]–which have their own complications [19, 20]. Although there is a push to use LLMs to perform automated evaluation, there is no consensus on the viability of this approach: arguments in favour [21, 22, 23, 24, 25] are as plentiful as arguments against [26, 27, 28, 29, 24, 30]. Finally, LLMs might memorise their training data [31, 32], including evaluation benchmarks [33]. Since the models are sensitive to the prompt’s phrasing [34, 35], the benchmarks turn into an arena to test prompts, rather than a holistic bar of performance.
Overall, these challenges impede the scientific community’s ability to transparently evaluate and understand LLMs, and to ultimately ensure their responsible use. They also raise questions about the validity and trustworthiness of some findings and claims related to this technology and its capabilities, and illustrate a problem endemic to computer science.
The goal of this work is to systematically and critically examine the scientific methodology employed in LLM research, and to quantify the extent to which these issues occur in the literature. [^2] Code and partial, anonymised data will be released in https://github.com/adewynter/awes_laws_and_flaws To our knowledge, ours is the first study to systematically evaluate LLM research based on the metrics the field recommends and requires. To do this, we evaluated 2,054 works that cited the peer-reviewed GPT-3 paper [36] (as opposed to the preprint) and the GPT-4 technical report [37], labelled based on criteria as shown in Table 1. Most criteria are part of the reproducibility checklist for premier conferences [^3] See, e.g. AAAI [38], NeurIPS [39] and Carpuat et al. [40] for ACL. and considered good practices in the field. Some criteria are not part of the checklists: the type of evaluator used is because of the aforementioned lack of consensus around the reliability of LLMs as evaluators, added to their increasingly frequent use. Claims of super-human capabilities, reasoning abilities, or emergent behaviour relate to conclusions often drawn and reported outside of academic channels, and that could require closer examination–for example, it is known that more robust statistical tests prove that emergent abilities do not actually occur [41].
@@
Table 1: Criteria for our analysis. The first three (Research Features, Structural Features, Arguments Made) are the core subject of our work. Most labels are binary labels (yes or no); but Research Features also include a “not applicable” label (“na”), and evaluators is a set ({human, automatic, LLM, na}). The latter (Indicators) is meant to act as a filter, since our focus is research articles with LLMs as the centre of the work. The label “book” also includes surveys. We use () for criteria typical of leading conference checklists, () for claims usually requiring closer examination, and (x) for those recommended–but not implemented in checklists–as good research practices. See Appendix A for definitions.
2 A Review of LLM Literature
We split our results in four analyses: corpus composition (Section 2.1), composition over time (Section 2.2), the relationship between citations and criteria (Section 2.3), and yearly trends on relationships between citations and criteria (Section 2.4). The full breakdown of results is in Appendix C. Throughout this section, we use relevant papers to refer to these that did not score “na” in the criterion.
@@ 2.1 How Many Papers Did What?
Out of the 2,054 articles surveyed, 57%, or 1,164, claimed SOTA results. From these, 30% contained or addressed ethical considerations related to their research; 13% performed evaluations in languages other than English; and 39% did not include limitation sections about their experimentation (Figure 1). Only 25% of them included statistical tests to support their claims, a number close to the 23% found by van der Lee et al. [12] three years ago. This number is lower on average than the papers that do not claim SOTA (23% versus 25%).
In terms of criteria overlap (Table 9), from the papers claiming SOTA and emergent capabilities (22%) only 26% relied on statistical tests or had error breakdowns (14%). Articles typically included automatic metrics (LLM without: 2%; human without: 4%), and reliance on LLM evaluators alone was exceedingly rare. Claims of reasoning were often done with LLM evaluators and not human evaluators (35%); contrasting with claims that they cannot reason predominantly done with human evaluation alone (14%). The use of automated metrics is more common (+10% on average). LLM evaluators remained steady, thus showing that most of the papers claiming SOTA used this metric in some form. A full breakdown by evaluator type is in Table 9.
On a positive note, a considerable amount of the relevant papers claiming SOTA did report model versioning (73%), and open-sourced their work (68%), a number slightly higher than the one found by Arvan et al. [13], indicating a positive trend.
@@
Figure 1: Breakdown of the corpus for selected criteria, narrowed down by papers claiming SOTA results. While open-sourcing, declaration of experimental protocols (e.g., parameters) and limitations are relatively high (60%+), there is a lack of ethics sections (30%) and evaluations in languages other than English (13%).
Criterion Corpus (%) Stat. Tests Error Bdown. Randomness Human Eval. LLM Eval. Automatic Eval.
Ethics 30 30 21 11 39 17 85
Limitations 61 31 18 11 32 16 86
Negative 13 40 32 15 39 7 81
Error Bdown. 16 33 100 16 49 17 95
Emergence 39 26 14 13 26 12 83
Can Reason 27 24 14 14 28 15 87
Cannot Reason 6 39 24 17 37 11 76
Superhuman 5 40 27 18 45 11 91
Human Eval. 27 34 29 10 100 16 84
LLM Eval. 13 26 22 13 34 100 82
Automatic Eval. 86 26 18 11 26 12 100
Mean (SOTA) 100 25 16 10 27 13 86
Table 2: Overlap for selected criteria with respect to Research Features. Highlighted are the claims that fall below the average for papers claiming SOTA. Claims that the models present emergent behaviour or that they models can reason are rarely backed by strong experimental protocols: only 14% of the papers for either criteria present error breakdowns–below the corpus average or any of the other criteria present. The use of LLM evaluators typically comes paired with error breakdowns and statistical tests. However, papers with negative results eschew their use.
2.2 Did it Change Over Time?
LLM research is undergoing an explosive–almost inflationary–growth. In our corpus 7% of the papers were from 2021; contrasting with 22% and 25% for 2022 and 2023, and behind the 46% for the first half of 2024. 2020 had a negligible sample size, given that the peer-reviewed GPT-3 paper was released on 6 December 2020.
For our time-wise analysis (Figure 2) we only worked with relevant papers claiming SOTA from 2021 onwards. Between 2023 and 2024 we observed declines in the absolute percentage of several metrics, such as the presence of ethics disclaimers (-3%), open-sourcing (-9%), claims of emergence (-7%), and statistical tests (-5%). Of note is a remarkable increase in the claims that these models can reason (15%); as opposed to that they cannot (-3%), as well as claims of emergence (-2%). On the other hand, limitations sections, error breakdowns, dialect evaluation, and relying on human evaluation have remained steady ($\pm$1%). Finally, the use of LLM evaluators has undergone an uptick (15%), despite the lack of consensus of the field about the reliability of this metric.

### v3

@@ Abstract
We perform a critical examination of the scientific methodology behind contemporary large language model (LLM) research. For this we assess over 2,000 research works released between 2020 and 2024 based on criteria typical of what is considered good research (e.g. presence of statistical tests and reproducibility), and cross-validate it with arguments that are at the centre of controversy (e.g., claims of emergent behaviour). We find multiple trends, such as declines in ethics disclaimers, a rise of LLMs as evaluators, and an increase on claims of LLM reasoning abilities without leveraging human evaluation. We note that conference checklists are effective at curtailing some of these issues, but balancing velocity and rigour in research cannot solely rely on these. We tie all these findings to findings from recent meta-reviews and extend recommendations on how to address what does, does not, and should work in LLM research.
[…]
[innertopmargin=-3pt, innerbottommargin=-3pt, innerleftmargin=3pt, innerrightmargin=-2pt, skipabove=0pt, skipbelow=0pt, backgroundcolor=gray!10, linecolor=black, outerlinewidth=0.01pt, font=, ]promptbox
Awes, Laws, and Flaws From Today’s LLM Research
Adrian de Wynter Microsoft and the University of York adewynter@microsoft.com
@@ 1 Introduction
Large language models (LLMs) [^1] We use the term ‘LLM’ loosely to refer to generative text-to-text models with sizes of or above 1B parameters. are a powerful technology. They can follow instructions and output coherent, persuasive text. This has made them the centre of attention in academia, industry, and the media. Given the potential funding associated with these technologies, it should not be a surprise that news and research works are sometimes accompanied by bold claims about their capabilities.
It has been said that the focus of AI research is the approaches, rather than the results. That is, less importance is allocated to issues around experimental protocols when compared to other fields Burnell et al. (2023). For example, papers sometimes lack sufficient details for independent verification Gehrmann et al. (2023); Hullman et al. (2022); or report aggregate performances (e.g., accuracy) without providing detailed protocols or error breakdowns Bouthillier et al. (2021); Gehrmann et al. (2023); Burnell et al. (2023); Liesenfeld et al. (2023); Hullman et al. (2022). This is common in health sciences McDermott et al. (2021), security Olszewski et al. (2023), and recommender systems Cremonesi and Jannach (2021)–all areas where LLMs are increasingly being applied.
@@
We argue, however, that in LLM-related research there has been a shift towards result-driven experimentation, partly due to their availability and capabilities; but also due to scrutiny from the media, funding sources, and the field. In theory, this should mean that the scientific community ought to allocate the same relevance to experimental protocols and good research practices as other fields.
This is easier said than done, since LLMs may be closed source, expensive to train, and/or limited to a versioned API call. Solving a problem could be done one single prompt, and their ability to generate text reduces the time taken to write and produce papers. A combination of these could lead to a rapidly-increasing volume of scientific articles with strong claims and lacklustre experimental practices. However, LLMs as a technology are not the only factor: evaluating these models is notoriously difficult, researchers are under pressure to keep up, and peer-reviewing systems are usually overloaded. Still, all of this impedes the scientific community’s ability to transparently evaluate and understand LLMs, and to ensure their responsible use. It also poses questions about the validity and trustworthiness of some findings and claims, especially around LLM capabilities.
Our work is a statistically motivated critique, meant to systematically and critically examine the scientific methodology employed in LLM research, and to quantify the extent to which these issues occur in the literature. [^2] Code and partial, anonymised data will be released in https://github.com/adewynter/awes_laws_and_flaws We look at what does, does not, and should work in LLM research, and, based on that, extend recommendations to the scientific community at large to retain velocity without sacrificing innovation or good research practices.
1.1 Contributions
We evaluate 2000+ scientific articles that have an LLM as the focus of study, and based on the presence of a set of criteria. Most of these come from reproducibility checklists for premier conferences. [^3] See, e.g. AAAI (2024); NeurIPS (2024) and Carpuat et al. (2024). The rest are chosen based on claims at the centre of controversy, such as using LLMs as evaluators, or assertions of emergent behaviour.
Our analysis shows various trends, such as a decreased emphasis on ethics or appropriate research protocols (accounting for versioning, declaring call parameters, etc.), and a rise on the use of LLMs as evaluators. There is also an increase of works in non-English languages and a steady number of limitation disclaimers.
The takeaways of this study, however, are that the field appears to be increasingly rushing on producing papers lacking rigour experimentally and ethically; while certain conference mechanisms, such as ACL’s enforcement of a limitations section, appear to be effective at curtailing some of these trends. This is important, as plenty of papers rely on recency and claims of SOTA for novelty–namely, citing non-peer reviewed sources. While not a problem per se, the robustness of the sources, and hence the paper’s arguments and results, could be put into question if they have not been carefully validated or used reliable experimental protocols. Take, for example, claims of emergent behaviour in LLMs Wei et al. (2022); Brown et al. (2020) and the fact that better statistical methods make them ‘evaporate’ Schaeffer et al. (2023).
It is difficult to maintain selectivity, novelty, and speed. We then suggest a middle ground where we educate ourselves and the community at large on what constitutes a paper with robust experimental practices. This middle ground comes as a series of recommendations to the field at large on how to maintain rigour in research, all without sacrificing velocity in research and innovation.
2 Background
Some known challenges to AI research are particularly salient in LLM literature, although might require reframing. In this section we describe and tie them to the criteria and focus of our analysis.
@@ 2 Background
Reproducibility: LLM-related reviews have found that the most downloaded models in Huggingface do not consistently provide the same amount of information in their documentation Liang et al. (2024); and an analysis on the experimental protocols of over 40 chat-based LLMs (e.g., availability of data, weights, licences, etc) found that, while many projects claimed some level of open-sourcing, documentation was ‘exceedingly rare’ Liesenfeld et al. (2023). Beyond that, reproducibility itself is tied to stochasticity and versioning, which in turn raises concerns about the results themselves, particularly around the fact that errors and biases are rarely reported or analysed Pang et al. (2025).
One core aspect of reproducibility, beyond protocol declaration, is open-sourcing. For LLM research in particular, this doesn’t readily apply to models behind APIs. While in the broader CS field open-sourcing has been trending upward, reproducibility remains difficult Arvan et al. (2022). Further, in some areas it has been found to be no statistically significant difference in the presence of artifacts (code, data) when the mechanisms for reproducibility were implemented Olszewski et al. (2023). This in turn suggests that, due to LLM stochasticity and system limitations, open-sourcing alone is preferrable, albeit not a cure-all: but protocol declaration, however, remains important.
Measurement is difficult in LLMs due to effectiveness and scalability of metrics. It has long been known that automated metrics like BLEU do not capture natural language generation well Liu et al. (2016); Novikova et al. (2017), and not correlate well with human judgements Reiter (2018); Gehrmann et al. (2023), which themselves have their own complications Clark et al. (2021); van der Lee et al. (2019). While it has been argued that metric performance is not as important as tasks and methods, and mostly drive improvement within the task Pramanick et al. (2023), benchmarks and leaderboards are the reigning way to define LLM success.
These have problems scaling, however, given how flexible LLMs are at multiple tasks. Although there is a push to use them as evaluators, there is no consensus on the viability of this approach: arguments in favour Chiang et al. (2024); Chiang and Lee (2023a); Liu et al. (2023); Wei et al. (2024); Zheng et al. (2023) are as plentiful as arguments against Doddapaneni et al. (2024); Stureborg et al. (2024); Chiang and Lee (2023b); de Wynter et al. (2025); Wei et al. (2024); Hada et al. (2024). One reason why this approach is often called unreliable is due to the fact that LLMs might memorise their training data Lee et al. (2023); de Wynter et al. (2023), including evaluation benchmarks Sainz et al. (2023); their results are very sensitive to the prompt’s phrasing Lu et al. (2022); Hida et al. (2024); and that, generally speaking, an NLG system’s performance, including evaluator systems, is strongly dependent on the choice of metric von Däniken et al. (2024); Gao et al. (2025). It is clear that carefully-chosen, diverse metrics and statistical tests are necessary for trustworthy results.
Claims: Many aspects of LLM research are related to their (or lack thereof) capabilities. For example, it is often said that LLMs present emergent abilities. This is usually defined as their ability to solve more complex problems in a way not predictable by, say, parameter size Wei et al. (2022); Brown et al. (2020). This definition is slightly vague, given that the problems or their measures of complexity are not actually defined. Another common claim associated with LLM research are claims of artificial general intelligence (AGI). Analogous to emergence, AGI claims often rely on disparate definitions and goals Blili-Hamelin et al. (2025), which in turn makes results incomparable. Even the term ‘reasoning’ is not always well-scoped (abductive? analogical? formal?; Huang and Chang 2023). The relationship of these claims to rigorous experimental protocols are central to our work. Aside, recall from our earlier arguments that some of these claims fall apart under closer scrutiny, namely under more robust statistical tests.
Ethics and Inclusion: LLMs are known to cause and propagate multiple harms, in addition to have very strong multilingual capabilities. It has then become an active focus area to ensure that LLMs are used in an ethical and inclusive manner for multiple audiences.
Note, however, that ethical concerns and evaluations are system-dependent: disclosures, protocols, and the concerns themselves are not universally tied to English-based language modelling, or even NLP. For example, alignment of models is usually carried out within a universal value system (e.g., ‘always be honest’), but this has been called out for being pragmatically inadequate Varshney et al. (2025). Likewise, evaluations under a single prompt tend to fail when working with culture-specific tasks Cheng and Hale (2025).
Add to that the fact that the risks called out by researchers may not be the same risks considered by laypeople Karamolegkou et al. (2024), or even other fields: only 2% of non-CS papers using LLMs call out ethical considerations Pramanick et al. (2024), perhaps due to less concern or familiarity with this area. Note that the focus of this work is LLM-centred research, so papers from fields other than NLP are also part of our analysis.
Hence, authors should have a responsibility to educate others on the impact of their technology, and be aware of what matters for broader audiences. Therefore, ethical and inclusive considerations should remain constantly present as part of any LLM-related work.
Recommendations: Studies and meta-reviews of AI often recommend good practices (e.g., reporting carbon footprints Henderson et al. (2020); reproducibility checklists and experimental protocols Pineau et al. (2021); Gundersen and Kjensmo (2018); van der Lee et al. (2021); self-contained artifacts Arvan et al. (2022); and metadata for corpora Gebru et al. (2021). These suggestions are often not heeded: Gehrmann et al. (2023) noted that, out of 66 articles from leading conferences, only between 15% and 35% of the recommendations were partially followed. The practices themselves might not even be sufficiently impactful Olszewski et al. (2023), given that other experimental protocols (e.g., sampling, initialisation, hyperparameters) may also impact reproducibility Bouthillier et al. (2021). In this work we are unable to directly address this lack of impact, although we do note in Section 7 where it is possible to have better momentum without (completely) enforcing checklists in conferences.
3 Methods
3.1 Corpus
Our corpus is comprised of works that cited the peer-reviewed GPT-3 paper Brown et al. (2020) and the GPT-4 technical report Open AI (2023). We make the assumption that the majority of the LLM literature references either of these articles. We examine this assumption more closely in Sections 6 and 9; and evaluate its longevity in a follow-up study done a year after our data cutoff (Appendix D).
We retrieved the top 1,000 papers sorted by citation numbers for both articles in Google Scholar [^4] https://scholar.google.com/ with Publish or Perish Harzing ; and the top 2,000 papers by citation number for GPT-3 in Scopus. [^5] https://www.scopus.com/ The disparity is due to Google Scholar indexing peer-reviewed works and preprints, and Scopus only indexing peer-reviewed articles. At the time of writing this paper, the GPT-4 technical report has yet to be peer-reviewed. All queries were ran for papers published or released up to 10 June 2024. Considerations around the representativeness of this corpus are in Section 9.
We used the arXiv API [^6] https://info.arxiv.org/help/api/index.html to retrieve the full paper, and parsed either the source into text. The final, deduplicated, unlabelled corpus is $3,914$ texts.
3.2 Evaluation Criteria
We labelled our corpus based on a set of criteria (labels), categorised in four groups: Research Features, Structural Features, Arguments Made, and Indicators. Groups and labels are in Table 1, with specific definitions–as prompts–in Appendix A.
Research Features, Structural Features, and Arguments Made are the core evaluation criteria used in the rest of our paper. Indicators is a filter for our corpus: we were only interested in research articles with an LLM as the subject of research.
@@ 3.2 Evaluation Criteria
Table 1: Criteria for our analysis. Research Features, Structural Features, and Arguments Made are the core subject of our work. Most labels are binary labels (yes/no); but Research Features also include ‘na’ (not applicable), and evaluators is a set ({human, automatic, LLM, na}). Indicators is a filter to only select LLM-centred work research articles. We use () for criteria from conference checklists and (x) for those recommended–but not implemented–as good research practices, and () for claims requiring closer examination. See Appendix A for definitions.
3.3 Labelling
Given the large volume of data and budget constraints, we were unable to perform a full human-based labelling work. Instead, we labelled the data with GPT-4 omni (version: gpt4-o-2024-05-13). To ensure reproducibility, we set the temperature to zero, the maximum output tokens to 256, and left other parameters as default. To improve accuracy we split in batches our labelling calls, totalling five different prompts Hada et al. (2024). All our calls were done through the Azure OpenAI API and the analysis done with a consumer-grade laptop.
To ensure trustworthiness of our results, we measured the model’s reliability by sampling 100 papers per criterion, and manually labelling them. We found the model’s results to be reliable to an average 91.91 $\pm$ 1.22% accuracy, with a 95% confidence interval. This number varies broadly across criteria. A full breakdown of reliability and analysis of performance is in Appendix B.
Regarding the label set, the model was instructed to return binary labels ({y, n}) for all criteria, except for most Research Features labels, which also included a relevance label ({na}). To determine the evaluators used in a given paper we used the set {human, LLM, automatic, na}; and for the type of text, {book, article, opinion}. For our topic analysis, we requested a primary subject for the paper in a few words, and then manually clustered them. Prior to our experiments, including topic analysis, we filtered the corpus by selecting all research articles with an LLM as the main subject of study. The final size of the data was $2,054$ papers.
4 A Review of LLM-Centred Literature
We provide four analyses: corpus composition (Section 4.1), composition over time (Section 4.2), the relationship between citations and criteria (Section 4.3), and yearly trends on relationships between citations and criteria (Section 4.4). The full breakdown of results is in Appendix C. Throughout this section, we use relevant papers to refer to these that did not score ‘na’ in the criterion discussed.
@@ 4.1 How Many Papers Did What?
We found that 57% of the articles claimed SOTA results. A third of these contained or addressed ethical considerations related to their research; 13% performed evaluations in languages other than English; and 39% did not include limitation sections about their experimentation (Figure 1). Only a quarter of them included statistical tests to support their claims–close to the 23% found by van der Lee et al. (2021). This is lower than for papers that do not claim SOTA. Further results are in Appendix C.1.
@@
Figure 1: Breakdown of the corpus for selected criteria, narrowed down by papers claiming SOTA results. While open-sourcing, declaration of experimental protocols, and limitations are relatively high (60%+); ethics sections (30%) and evaluations in languages other than English (13%) are comparatively low.
In terms of criteria overlap (Figure 2), from the papers claiming SOTA and emergent capabilities, only a quarter of them relied on statistical tests or had error breakdowns. The articles usually included automatic metrics, and reliance on LLM evaluators alone was exceedingly rare. Claims of LLM reasoning capabilities were often done with LLM evaluators and not human evaluators; conversely, claims that they cannot reason were done with human evaluation alone.
Figure 2: Percentage overlap for selected criteria with respect to Research Features. The claims whose percentage is below the average for papers claiming SOTA are marked with a down arrow.
Compared to the entire corpus, SOTA papers use fewer measures of statistical significance overall, and the use of automated metrics is more common than other types of measurements. That said, SOTA papers only using LLMs as evaluators were rare. Many relevant SOTA papers did report model versioning (73%), and open-sourced their work (68%), a number slightly higher than the one found by Arvan et al. (2022), indicating growth. A full breakdown of the results is in Appendix C.2.
4.2 What Changed Over Time?
In our corpus, 46% of the papers belonged to the first half of 2024, contrasting with 22% and 25% for 2022 and 2023. Given that our 2024 subset only comprised half of a year–but twice as many works as in 2022 or 2023–we may assume that this analysis will not incur any recency bias.